Could I get some help with this please? Thank you!!
Answer #1

If the ratio of volumes of the solution of [CoCl2(EtOH)2] and methanol is x: y and the total volume is 4 mL, then the volume that corresponds to the solution of [CoCl2(EtOH)2] = {x/(x + y)}*4 mL and the volume corresponds to methanol = {y/(x + y)}*4 mL

Ratio of volumes Volume of the solution of [CoCl2(EtOH)2] (mL) Volume of methanol (mL)
3: 1 (3/4)*4 = 3 1
4: 1 (4/5)*4 = 3.2 0.8
5: 1 (5/6)*4 = 3.333 0.667

and